20 cm 20 cm 20 cm KN 4 kN 35 mm

2014-T6 Aluminum

T1 = 22 degrees C. Then it is loaded, and temp changes to T2 = 0 degrees C
What is the change in length?
Answer #1

solution - Draw free Body Diagram properties of Aluminium 2014 - T6 Modulus of Elasticity (E)= 73. 1 G Pa Coefficient of ther= 12 x 103 L АЄ A = 0, + A2 + g = + ( 22 +30 +12 ) x 10 x 200 II x 35²x73-1x103 10 = +0.182 mm) 2 (tive sign indicate ? 2 exp